# Large Language Model Adaptation for Networking

### v1

@@ Abstract.
Many networking tasks now employ deep learning (DL) to solve complex prediction and system optimization problems. However, current design philosophy of DL-based algorithms entails intensive engineering overhead due to the manual design of deep neural networks (DNNs) for different networking tasks. Besides, DNNs tend to achieve poor generalization performance on unseen data distributions/environments.
Motivated by the recent success of large language models (LLMs), for the first time, this work studies the LLM adaptation for networking to explore a more sustainable design philosophy. With the massive pre-trained knowledge and powerful inference ability, LLM can serve as the foundation model, and is expected to achieve “one model for all” with even better performance and stronger generalization for various tasks. In this paper, we present NetLLM, the first LLM adaptation framework that efficiently adapts LLMs to solve networking problems. NetLLM addresses many practical challenges in LLM adaptation, from how to process task-specific information with LLMs, to how to improve the efficiency of answer generation and acquiring domain knowledge for networking. Across three networking-related use cases - viewport prediction (VP), adaptive bitrate streaming (ABR) and cluster job scheduling (CJS), we showcase the effectiveness of NetLLM in LLM adaptation for networking. Results show that the adapted LLM surpasses state-of-the-art algorithms by 10.1-36.6% for VP, 14.5-36.6% for ABR, 6.8-41.3% for CJS, and also achieves superior generalization performance.
@@
copyright: acmcopyrightcopyright: acmlicensedcopyright: rightsretainedcopyright: usgovcopyright: usgovmixedcopyright: cagovcopyright: cagovmixed
@@ 1.1. The Main Roadmap so far
Over the past decades, rule-based algorithms built on handcrafted control rules have played an important role in optimizing network systems (Arun and Balakrishnan, 2018; Grandl et al., 2014; Bai et al., 2015; Yin et al., 2015). For instance, Copa (Arun and Balakrishnan, 2018) adjusts sending rates for congestion control based on measured queueing delay, while PANDA (Li et al., 2014) switches video streaming bitrates based on heuristically estimated bandwidth. However, these algorithms heavily rely on rule engineering, which involves intensive human efforts to devise, implement and validate the control rules for network optimization (Chen et al., 2018a; Meng et al., 2020; Mao et al., 2019, 2017).
@@
  • High model engineering costs. The focus of learning-based algorithms has shifted from rule engineering to model engineering. Their success is heavily dependent on the manual engineering of DNN models for specific networking tasks, which, however, can be labor-intensive due to the complex structures of DNNs (Meng et al., 2020). To make things worse, the diversity of networking tasks also prevents sharing the same DNN model across different tasks. This necessitates designing specialized DNNs for different tasks (i.e., one model for one task), thus further increasing the engineering costs.
  • Low generalization. DNNs trained on specific data distributions/environments may struggle to perform well or even worse than conventional rule-based algorithms on unseen data distributions/environments (Yan et al., 2020). For example, an ABR model trained on smooth network conditions often achieves poor performance on network environments with dynamic bandwidth fluctuations (Kan et al., 2022). The lack of generalization can ultimately hinder the widespread deployment of learning-based algorithms in practice (Xia et al., 2022; Yen et al., 2023).
@@ 1.2. New Opportunities and Challenges
Utilizing a single generic model for different tasks has been recognized as a significant approach to mitigate the costs of handcrafting specialized DNNs for each task and enhance generalization (Reed et al., 2022). This is exemplified by the recent popular large language models (LLMs) such as ChatGPT (OpenAI, 2024), Flacon (Penedo et al., 2023) and Llama2 (Touvron et al., 2023) in the field of natural language processing (NLP). With billions of parameters pre-trained on massive data to absorb extensive knowledge, LLMs have demonstrated extraordinary capacity in conversations, reasoning and text generation in NLP (Wei et al., 2022). What’s more, their knowledge in pattern mining, problem solving, and generalization to unseen conditions has proven to be transferable to other domains (Zitkovich et al., 2023), including robotics (Driess et al., 2023), protein structure prediction (Lin et al., 2023) and chip design (Liu et al., 2023a). With these inspiring outcomes, we believe LLMs can serve as the foundation models for networking, as many networking tasks can also benefit from their extensive knowledge. For instance, in the context of ABR and congestion control, the pattern mining ability of LLMs can be utilized to mine complex changing patterns of bandwidth and delay for better decisions on bitrates and sending rates. Moreover, their generalization capability can be harnessed to generalize across diverse network environments. Therefore, we envision LLM as the key to achieve one model for all tasks, with little handcraft costs and strong generalization. We try to answer the following key question: can we embrace the era of LLMs, and easily adapt LLMs to solve various networking tasks efficiently and effectively?
@@
Unfortunately, as revealed by our analysis in $\S$3, the adaptation of LLM for networking faces the following challenges.
@@
  • Large input modality gap. In networking tasks, various information observed from the system is collected as inputs for networking algorithms (e.g., network throughputs and playback buffer length for ABR algorithms (Mao et al., 2017)). However, the modalities of these inputs differ significantly from plain text, i.e., the native input modality supported by LLMs [^1] Although multimodal LLMs have emerged recently (e.g., GPT4 (OpenAI, 2023)), their supported input modalities are still limited (mainly vision or audio), hindering directly applying them to process networking task inputs. (Touvron et al., 2023; Zhang et al., 2022a). This discrepancy prevents LLMs to effectively process the input information of networking tasks.
@@
  • Inefficiency of answer generation. LLM generates answers using a language modeling (LM) head to predict words (tokens) one by one (see Figure 1) (Kwon et al., 2023). While this approach is well-suited in NLP, it presents several drawbacks in the networking domain. First, LLM is prone to hallucination due to the inherent uncertainty of token prediction (Ji et al., 2023; Li et al., 2023a). Their generated answers for networking tasks may seem correct but physically invalid (e.g., a nonexistent bitrate for video download in ABR), which can eventually impair the reliability of network systems. Second, since tokens are predicted one at a time, LLM often requires multiple inferences to generate a complete answer, thus incurring high answer generation latency (i.e., the time to generate a complete answer). Consequently, LLM may fail to quickly generate answers to respond to the changes in network systems (e.g., switching to a lower bitrate when network bandwidth becomes scarce).
@@
  • High adaptation costs. The large domain gap between networking and NLP necessitates fine-tuning LLM to acquire domain-specific knowledge for effective adaptation. However, the adaptation costs can be prohibitively high, especially when fine-tuning LLM for decision-making tasks (e.g., ABR (Kan et al., 2022) and CJS (Mao et al., 2019)) where RL is employed to solve the system optimization problems. Specifically, RL-based decision-making tasks require the active interaction between LLM and environments (e.g., network environments with varying bandwidth or workload patterns) to collect experiences for performance optimization (Mao et al., 2019; Xia et al., 2022). Due to the large parameter size of LLM, the interaction process can be excessively time-consuming, introducing a large amount of additional training time. What’s worse, the adaptation costs can further increase if fine-tuning the full parameters of LLM, which is known to be resource-intensive (Lin et al., 2023; Liu et al., 2023a).
@@ 1.3. Design and Contributions
This work presents NetLLM, the first unified LLM adaptation framework that efficiently adapts LLMs to solve networking tasks. Specifically, NetLLM comprises the following three core designs to overcome the aforementioned challenges:
@@
  • Multimodal encoder. NetLLM designs an efficient multimodal encoder at the input side of LLM to effectively process the multimodal input information of networking tasks. At the core of this module, NetLLM uses modality-specific feature encoders to extract features from raw inputs. It then leverages trainable linear layers to project these features into token-like embedding vectors, which can be directly fed into the LLM for effective processing.
@@
  • Networking head. To enable efficient answer generation, NetLLM removes the default LM head used by LLM for token prediction. Instead, it introduces various networking heads at the output side of LLM to generate answers for specific networking tasks. Each networking head is essentially a lightweight trainable projector that maps the output features of LLM directly into task-specific answers. This enables LLM to generate a valid answer in a single inference, thus ensuring the reliability of LLM for networking and significantly reducing generation latency.
@@
  • Data-driven low-rank networking adaptation (DD-LRNA). To reduce the adaptation costs, NetLLM designs a DD-LRNA scheme for LLM to efficiently acquire domain knowledge for networking. Specifically, DD-LRNA incorporates a data-driven adaptation pipeline to adapt LLM for both prediction and decision-making tasks. In particular, for decision-making tasks, it employs the efficient data-driven RL technique (Yen et al., 2023; Agarwal et al., 2020; Prudencio et al., 2023) to eliminate the time-consuming interaction between LLM and environments. It collects an experience pool as training dataset with existing methods and fine-tunes LLM over such dataset in the data-driven manner. Besides, inspired by the advanced parameter-efficient fine-tune technique (Ding et al., 2023; Fu et al., 2023b), DD-LRNA introduces a set of additional trainable low-rank matrices for LLM to learn networking knowledge. Since the low-rank matrices only account for 0.31% of the total parameters, the fine-tune costs are greatly reduced, with the reduction of 60.9% GPU memory and 15.1% training time.
@@
NetLLM has some important properties: i) Compatibility: it is independent of specific LLMs or networking tasks. That said, different LLMs can be easily adapted with this unified framework to solve a variety of networking tasks, by using the multimodal encoders to process various task inputs, generating task-specific answers with networking heads, and acquiring domain knowledge with the DD-LRNA scheme. ii) Reliability: it addresses the hallucination issue and ensures the LLM generated answers to be always valid. iii) Efficiency: the DD-LRNA scheme significantly reduces the costs of fine-tuning LLM to learn domain knowledge and the networking heads also reduce the answer generation latency.
@@
We have implemented NetLLM [^2] The codes of NetLLM will be publicly available after acceptance. for three networking tasks: viewport prediction (VP) (Rondón et al., 2022) for immersive video streaming, adaptive bitrate streaming (ABR) (Xia et al., 2022), and cluster job scheduling (CJS) (Mao et al., 2019). We believe these representative tasks cover the main input modalities of networking problems and span from prediction tasks to decision-making tasks ($\S$3). Through extensive trace-driven simulation and real-world tests, we demonstrate the effectiveness of NetLLM in LLM adaptation for networking. We showcase that across the three use cases, LLM adapted by NetLLM significantly outperforms state of the arts with performance improvements of 10.1-36.6% for VP, 14.5-36.6% for ABR, 6.8-41.3% for CJS. Besides, our empirical test results also show that we can efficiently utilize the extensive knowledge of LLM with the proposed NetLLM framework to achieve stronger generalization on unseen testing environments.
@@
  • We then design NetLLM, the first unified LLM adaptation framework for networking that incorporates a multimodal encoder module to encode multimodal task inputs, networking head module to directly generate answers and a DD-LRNA scheme to reduce the adaptation costs ($\S$4).
  • We extensively evaluate NetLLM across three networking tasks. We showcase that LLM adapted by our framework can significantly surpass state-of-the-art algorithms and achieve superior generalization performance. We also provide in-depth understanding of NetLLM and shed light on some practical concerns when using NetLLM to adapt LLMs for networking ($\S$5).
Figure 1. Illustration of LLM answer generation mechanism.
@@ 2.1. Learning-Based Networking Algorithms
The major limitations of learning-based algorithms are two-fold. First, they require manual design of specialized DNN models for different networking tasks, thus entailing high model engineering overhead (Meng et al., 2020). Second, they are prone to generalization issues, as DNNs may achieve poor performance on unseen data distributions/environments (Kan et al., 2022).

### v3

# NetLLM: Adapting Large Language Models for Networking

@@ Abstract.
Many networking tasks now employ deep learning (DL) to solve complex prediction and optimization problems. However, current design philosophy of DL-based algorithms entails intensive engineering overhead due to the manual design of deep neural networks (DNNs) for different networking tasks. Besides, DNNs tend to achieve poor generalization performance on unseen data distributions/environments.
Motivated by the recent success of large language models (LLMs), this work studies the LLM adaptation for networking to explore a more sustainable design philosophy. With the powerful pre-trained knowledge, the LLM is promising to serve as the foundation model to achieve “one model for all tasks” with even better performance and stronger generalization. In pursuit of this vision, we present NetLLM, the first framework that provides a coherent design to harness the powerful capabilities of LLMs with low efforts to solve networking problems. Specifically, NetLLM empowers the LLM to effectively process multimodal data in networking and efficiently generate task-specific answers. Besides, NetLLM drastically reduces the costs of fine-tuning the LLM to acquire domain knowledge for networking. Across three networking-related use cases - viewport prediction, adaptive bitrate streaming and cluster job scheduling, we showcase that the NetLLM-adapted LLM significantly outperforms state-of-the-art algorithms.
@@
copyright: acmlicensedjournalyear: 2024doi: https://doi.org/10.1145/3651890.3672268journalyear: 2024conference: ACM SIGCOMM 2024 Conference; August 4–8, 2024; Sydney, NSW, Australiabooktitle: ACM SIGCOMM 2024 Conference (ACM SIGCOMM ’24), August 4–8, 2024, Sydney, NSW, Australiadoi: 10.1145/3651890.3672268isbn: 979-8-4007-0614-1/24/08isbn: 979-8-4007-0614-1/24/08ccs: Networks Application layer protocolsccs: Computing methodologies Supervised learningccs: Computing methodologies Reinforcement learning
@@ 1.1. The Main Roadmap so far
Over the past decades, rule-based algorithms built on handcrafted control rules have played an important role in optimizing network systems (Arun and Balakrishnan, 2018; Grandl et al., 2014; Bai et al., 2015; Yin et al., 2015). For instance, Copa (Arun and Balakrishnan, 2018) adjusts sending rates for congestion control based on measured queueing delay, while PANDA (Li et al., 2014) switches video streaming bitrates based on heuristically estimated bandwidth. However, these algorithms heavily rely on rule engineering, which involves intensive human efforts to devise, implement and validate the control rules for network optimization (Chen et al., 2018; Meng et al., 2020; Mao et al., 2019, 2017).
@@
  • High model engineering costs. It has been shown that the design of DNN architecture is crucial to the final performance (Miikkulainen et al., 2024; Ren et al., 2021). Therefore, the focus of learning-based algorithms has shifted from rule engineering to model engineering. Their success is heavily dependent on engineering DNN models for the target networking tasks, which, however, can be difficult and labor-intensive due to the complex structures of DNNs (Meng et al., 2020). To make things worse, the diversity of networking tasks also prevents sharing the same DNN model across different tasks. This necessitates designing specialized DNNs for different tasks (i.e., one model for one task), thus further increasing the engineering costs. Although some recent works attempt to introduce structured Transformer (Vaswani et al., 2017) into model design, they still necessitate manual tuning of the Transformer architecture (e.g., the number of attention blocks and attention heads) (Wu et al., 2023), or even the design of specialized tokenization scheme (Guthula et al., 2023; Meng et al., 2023) and attention mechanism (Liu et al., 2023e; Guthula et al., 2023), thus leading to high engineering costs.
  • Low generalization. DNNs trained on specific data distributions/environments may struggle to perform well or even worse than conventional rule-based algorithms on unseen data distributions/environments (Yan et al., 2020). For example, an ABR model trained on smooth network conditions often achieves poor performance on network environments with dynamic bandwidth fluctuations (Kan et al., 2022). The lack of generalization can ultimately hinder the widespread deployment of learning-based algorithms in practice (Xia et al., 2022; Yen et al., 2023), as network operators will suspect their superiority over the incumbent rule-based algorithms in production environments.
@@ 1.2. New Opportunities and Challenges
Utilizing a single generic model for different tasks has been recognized as a significant approach to mitigate the costs of handcrafting specialized DNNs for each task and enhance generalization (Reed et al., 2022). This is exemplified by the recent popular large language models (LLMs) such as ChatGPT (OpenAI, 2024), Falcon (Penedo et al., 2023) and Llama2 (Touvron et al., 2023) in the field of natural language processing (NLP). With billions of parameters pre-trained on massive data to absorb extensive knowledge, LLMs have demonstrated extraordinary capacity in conversations, reasoning and text generation in NLP (Wei et al., 2022). What’s more, they also exhibit emergent abilities that were not explicitly programmed into them during pre-training, such as planning, pattern mining, problem solving and generalization to unseen conditions (Xi et al., 2023; Zitkovich et al., 2023). These abilities have been proven to be transferable to other domains, including robotics (Driess et al., 2023), chip design (Liu et al., 2023a), and protein structure prediction (Lin et al., 2023). For instance, researchers have shown that LLMs can generate goal-oriented plans for robotic control, adjusting plans in response to environment changes, and generalize to previously unseen operating environments (Zitkovich et al., 2023).
Motivated by these inspiring outcomes, we believe that LLMs can serve as foundation models for networking, as many networking tasks can also benefit from their emergent abilities. In the context of ABR, for example, the planning ability of LLMs can be utilized for better bitrate decisions to optimize the video streaming sessions based on the changing network conditions. Furthermore, their generalization capability can be harnessed to generalize across diverse network environments. Therefore, we envision LLM as the key to achieving one model for all tasks, with little handcraft costs and strong generalization. We try to answer the following key question: can we embrace the era of LLMs and adapt LLMs to solve various networking tasks efficiently and effectively?
@@
Unfortunately, as revealed by our analysis in $\S$3, the adaptation of LLMs for networking faces the following challenges.
@@
  • Large input modality gap. In networking tasks, various information observed from the system is collected as inputs for networking algorithms. However, the modalities of these inputs differ significantly from plain text, i.e., the native input modality supported by LLMs [^1] Although multimodal LLMs have emerged recently (e.g., GPT4 (OpenAI, 2023)), their supported input modalities are still limited (mainly vision or audio), hindering directly applying them to process networking task inputs. (Touvron et al., 2023; Zhang et al., 2022a). For example, in ABR, network throughputs and delay are often collected for bitrate decision (Mao et al., 2017), which exhibit unique time-varying data patterns typically not found in natural text. This discrepancy prevents LLMs from effectively processing the input information of networking tasks.
@@
  • Inefficiency of answer generation. LLMs generate answers using a language modeling (LM) head to predict words (tokens) one by one (see Figure 1) (Kwon et al., 2023). While this approach is well-suited in NLP, it presents several drawbacks in the networking domain. First, LLMs are prone to hallucination due to the inherent uncertainty of token prediction (Ji et al., 2023; Li et al., 2023a). Their generated answers for networking tasks may seem correct but physically invalid (e.g., a nonexistent bitrate for video download in ABR), which can eventually impair the reliability of network systems. Second, since tokens are predicted one at a time, LLMs often require multiple inferences to generate a complete answer, thus incurring high answer generation latency (i.e., the time to generate a complete answer). Consequently, LLMs may fail to quickly generate answers to respond to the system changes (e.g., switching to a lower bitrate when network bandwidth becomes scarce).
@@
  • High adaptation costs. The large domain gap between networking and NLP necessitates fine-tuning LLMs to acquire domain-specific knowledge for effective adaptation. However, the adaptation costs can be prohibitively high, especially when fine-tuning LLMs for decision-making tasks (e.g., ABR (Kan et al., 2022) and CJS (Mao et al., 2019)) where RL is employed to solve the system optimization problems. Specifically, RL-based decision-making tasks require the active interaction between LLMs and environments (e.g., network environments with varying bandwidth or workload patterns) to collect experiences for performance optimization (Mao et al., 2019; Xia et al., 2022). Due to the large parameter size of LLMs, the interaction process can be excessively time-consuming, introducing a large amount of additional training time. What’s worse, the adaptation costs can further increase if fine-tuning the full parameters of LLMs, which is known to be resource-intensive (Lin et al., 2023; Liu et al., 2023a).
@@ 1.3. Design and Contributions
To overcome the aforementioned challenges, this work proposes NetLLM, the first framework that efficiently adapts LLMs for networking. NetLLM stands out as the first to take LLMs the extra mile in context of networking and provides a coherent design to utilize their powerful capabilities to generalize to various tasks with low efforts. It enables the efficient utilization of a single LLM (e.g., Llama2 (Touvron et al., 2023)) as the foundation model without any modifications to tackle a wide range of networking tasks and meanwhile achieves enhanced performance. To accomplish this, NetLLM designs a multimodal encoder to enable the LLM to process multimodal data in networking and implements a networking head to improve the efficiency of answer generation. Furthermore, while fine-tuning the LLM and these two modules on the target task to acquire domain knowledge is necessary, NetLLM designs an efficient data-driven low-rank networking adaptation (DD-LRNA) scheme to drastically reduce the fine-tuning costs. Specifically, NetLLM incorporates the following three core design modules to efficiently adapt LLMs for networking.
@@
  • Multimodal encoder. NetLLM designs an efficient multimodal encoder at the input side of the LLM to effectively process the multimodal input information of networking tasks. The goal of this module is to automatically project task inputs to the same feature space as language tokens, so that the LLM can understand and utilize these inputs for task solving. To achieve this, it first uses modality-specific feature encoders to extract features from raw inputs of various modalities involved in networking. It then leverages trainable layers to project these features into token-like embedding vectors, which can be directly fed into the LLM for effective processing.
@@
  • Networking head. To enable efficient answer generation, NetLLM removes the default LM head used by the LLM for token prediction. Instead, it introduces various networking heads at the output side of the LLM to generate answers for specific networking tasks. Each networking head is essentially a lightweight trainable projector that maps the output features of the LLM directly into task-specific answers. In other words, they eliminate token prediction and allow direct answer generation from the valid range of possible answers (e.g., selecting a bitrate from candidate options in ABR). This enables the LLM to generate a valid answer in a single inference, thus ensuring its reliability for networking and significantly reducing generation latency.
@@
  • DD-LRNA. To reduce the fine-tuning costs, NetLLM designs a DD-LRNA scheme for the LLM to efficiently acquire domain knowledge for networking. Specifically, DD-LRNA incorporates a data-driven adaptation pipeline to adapt the LLM for both prediction and decision-making tasks. In particular, for decision-making tasks, it employs the efficient data-driven RL technique (Yen et al., 2023; Agarwal et al., 2020; Prudencio et al., 2023) to eliminate the time-consuming interaction between the LLM and environments. It collects an experience pool as training dataset with existing networking algorithms and fine-tunes the LLM over such dataset in the data-driven manner. Besides, inspired by the advanced parameter-efficient fine-tune technique (Ding et al., 2023; Fu et al., 2023b), DD-LRNA introduces a set of additional trainable low-rank matrices for the LLM to learn networking knowledge. Since the low-rank matrices only account for 0.31% of the total parameters, the fine-tune costs are greatly reduced, with the reduction of 60.9% GPU memory and 15.1% training time.
@@
NetLLM has the following important properties. i) Compatibility: It is a generic framework that can be applied to adapt different LLMs for various networking tasks. ii) Reliability: It addresses the hallucination issue and ensures the LLM generated answers to be always valid. iii) Efficiency: The DD-LRNA scheme significantly reduces the costs of fine-tuning LLM to learn domain knowledge and the networking heads also reduce the answer generation latency. iv) Transferability: NetLLM offers an effective solution for utilizing LLMs to solve prediction and decision-making tasks at low costs. These two types of tasks, in fact, span across diverse domains such as medicine (Clusmann et al., 2023), finance (Li et al., 2023c), and telecommunication (Zhou et al., 2024). Hence, despite its original focus on networking, NetLLM holds the potential to be applied and transferred to other fields.
@@
We have implemented NetLLM [^2] The codes available at: https://github.com/duowuyms/NetLLM. for three networking tasks: viewport prediction (VP) (Rondón et al., 2022) for immersive video streaming, adaptive bitrate streaming (ABR) (Xia et al., 2022), and cluster job scheduling (CJS) (Mao et al., 2019). We believe these representative tasks cover the main input modalities of networking problems and span from prediction tasks to decision-making tasks ($\S$3). Through extensive trace-driven simulation and real-world tests, we demonstrate the effectiveness of NetLLM in LLM adaptation for networking. We showcase that across the three use cases, the NetLLM-adapted LLM significantly outperforms state of the arts with performance improvements of 10.1-36.6% for VP, 14.5-36.6% for ABR, 6.8-41.3% for CJS, in terms of their respective performance metrics. Besides, our empirical test results also show that we can efficiently utilize the extensive knowledge of the LLM with our proposed NetLLM framework to achieve stronger generalization on unseen testing environments.
@@
  • We then design NetLLM, the first LLM adaptation framework for networking that incorporates a multimodal encoder module to encode multimodal task inputs, networking head module to directly generate answers and a DD-LRNA scheme to reduce the adaptation costs ($\S$4).
  • We extensively evaluate NetLLM across three networking tasks. We showcase that the LLM adapted by our framework can significantly surpass state-of-the-art algorithms and achieve superior generalization performance. We also conduct a deep dive into NetLLM to provide an in-depth understanding of it ($\S$5).
@@ 2.1. Learning-Based Networking Algorithms
The major limitations of learning-based algorithms are two-fold. First, they require designing specialized DNN models for different networking tasks, thus entailing high model engineering overhead (Meng et al., 2020). Second, they are prone to generalization issues, as DNNs may achieve poor performance on unseen data distributions/environments (Kan et al., 2022). These problems, in fact, are not unique to networking and have been successfully addressed in other domains (e.g., NLP (Kojima et al., 2022; Brown et al., 2020) and robotics (Zitkovich et al., 2023; Driess et al., 2023)) by introducing LLMs as the foundation models to solve various downstream tasks. Hence, this work takes a pioneering step by systematically introducing LLMs into networking to address these problems.